OpenStudy (anonymous):

G and H are mutually exclusive events. P (G) = 0.5; P (H) = 0.3 b. Find: P(H OR G).

OpenStudy (anonymous):

Always P(H or G) = P(G) + P(H) - P(GnH) Using Venn Diagram this equality is very visible. When Events are ME the intersection is ZERO because they can't happen simultaneously. so P(H or G) = P(H) + P(G) = 0.8

OpenStudy (anonymous):